parents set up a trust fund for their children last year. part of it is earning 7% simple interest per year while the rest of is is earning 6% simple interest per year. They placed $6000 in the trust fund. in one year the trust fund has earned $380. how much did they invest at each interest rate?
Answer by macston(5194)

.
x=amount at 6%; y=amount at 7%
x+y=$6000
x=$6000-y
.
0.06x+0.07y=$380 Substitute for x from above.
0.06($6000-y)+0.07y=$380
$360-0.06y+0.07y=$380
0.01y=$20
y=$2000 ANSWER 1: $2000 was invested at 7%.
.
x=$6000-y=$6000-$2000=$4000 ANSWER 2: $4000 was invested at 6%.
.
CHECK:
0.06x+0.07y=$380
0.06($4000)+0.07($2000)=$380
$240+$140=$380
$380=$380